OpenStudy (anonymous):

A Punter Kicks A Football . Its Height (h) in meters ,t seconds after the kicks is given by the equation :h=-4.9t^2 + 18.24t +0.8 . The Height Of An Approaching Blockers Hands Is Modeled By Equation : h=-1.43t + 4.26 , using the same time . Can The Blocker Knock Down The Punt ? If So , At What Time Does It Happen ?

OpenStudy (jtvatsim):

This one is a little tricky, but I think the plan should be: 1) Set the equations equal to each other (we want to see if the football touches the hands). 2) We will probably need to use the quadratic formula to solve. Do you want to give that a try? I'll help guide you through it.

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Alright, let's start with the first step. We will set the equations equal to start: -4.9t^2 + 18.24t + 0.8 = -1.43t + 4.26

OpenStudy (jtvatsim):

Now, we should probably try to set the equation equal to zero. I'd suggest moving the -1.43t + 4.26 to the other side, like this:

OpenStudy (jtvatsim):

-4.9t^2 + 19.67t +5.06 = 0 Do you see how I got that?

OpenStudy (anonymous):

Umm I Think So

OpenStudy (anonymous):

Adding Correcct? 1.43 + 4.26 TO BOTH SIDES

OpenStudy (jtvatsim):

Oh wait, lol. Good catch. I was supposed to add the 1.43t, but subtract the 4.26. Sorry about that. :)

OpenStudy (jtvatsim):

-4.9t^2 + 19.67t - 3.46 = 0 There... that's better. :P

OpenStudy (jtvatsim):

That should make more sense now. OK so far?

OpenStudy (anonymous):

YES

OpenStudy (jtvatsim):

Excellent! Alright, now for the "ugly" part. We have really nasty decimal numbers here, so we should probably just use the quadratic formula. Do you remember the formula or should we look it up?

OpenStudy (anonymous):

i Remember

OpenStudy (jtvatsim):

Alright, so I'll trust you and spare you the pain of typing it all out. I will go ahead and plug the numbers into the formula, and we can compare answers after. :)

OpenStudy (jtvatsim):

What we can definitely say is that: a = -4.9 b = 19.67 and c = -3.46 for when we plug in.

OpenStudy (jtvatsim):

\[\frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-19.67\pm\sqrt{19.67^2-4(-4.9)(-3.46)}}{2\cdot -4.9}\]

OpenStudy (anonymous):

Exactly Correct

OpenStudy (jtvatsim):

Perfect! Now to simplify that horrid looking nightmare. I'll use a calculator now...

OpenStudy (jtvatsim):

One answer is t = 0.184... the other....

OpenStudy (jtvatsim):

t = 3.830

OpenStudy (anonymous):

Yes

OpenStudy (jtvatsim):

OK, but now what? We have two possibilities either the blocker grabs the ball at t = 0.184 or at t = 3.830... we need a way to figure out which one.

OpenStudy (jtvatsim):

But we should at least celebrate our progress so far. YAY!!! :)

OpenStudy (anonymous):

Yay !!!!!!!!!!!!! Lol

OpenStudy (jtvatsim):

OK, let's do a sanity check with our answers. A useful algebra trick is to plug answers back into the original equations to see if they make sense.

OpenStudy (jtvatsim):

Let's not use the -4.9t^2 + ... mess, instead I'll check the other equation: -1.43t + 4.26

OpenStudy (anonymous):

Right Thats Just Makes Us Go To Far.

OpenStudy (jtvatsim):

OK, plugging in t = 0.184, I see h = 3.99 But t = 3.830, I see h = -1.22

OpenStudy (jtvatsim):

Well, there's a big problem with the second answer. Do you see it? The blocker's hands are a "negative height," in other words he blocks the ball underground. Unless he is a gopher, I think we can conclude that we should go with the first answer t = 0.184

OpenStudy (anonymous):

No I Didnt See It , i Just Re-Read The Problem WOW

OpenStudy (anonymous):

Almost Missed It

OpenStudy (jtvatsim):

Yep, as I mentioned when we first started, this one was a little sneaky. But hopefully that makes sense overall. Any final questions on this one? :)

OpenStudy (anonymous):

Yes .

OpenStudy (anonymous):

Well I Have A Question Mind Helping With Another Onne ?

OpenStudy (jtvatsim):

Go ahead and close this one and post a new one, so it doesn't get too cluttered. See you there! :)

OpenStudy (anonymous):

Ok